6.66 points Given the following Number purchasedper unitTotal Cost January 1 inventory April 1 June 1 November 1 $3$ 105 330 315 400 $ 1.150 35 45 50 185 a. Calculate the cost of ending inventory using the LIFO (ending inventory shows 58 units) Cost of ending inventory b. Calculate the cost of goods sold using the LIFO (ending inventory shows 58 units) Cost of goods sold References eBook & Resources Worksheet Difficulity 1 Easy 18-01 (2) Calculat ending inventony nd pest of goods inventory nethod
Answer #1

a) Cost of ending Inventory using LIFO = 35*3+21*6 = 105 + 126 = 231

b) Items Sold = 185 - 56 = 129
Cost of Goods Sold = 50*8 + 45*7 + 34*6 = 919

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well